TUNNELING SPLITTINGS FOR THE NOMINAL OO STRETCHING AND OTHER VIBRATIONS OF TROPOLONE ISOTOPOMERS OBSERVED IN THE LOW FREQUENCY INFRARED SPECTRUM

\underline{RICHARD L. REDINGTON} and THERESA E. REDINGTON, Department of Chemistry and Biochemistry, Texas Tech University, Lubbock, TX 79409; ROBERT L. SAMS, Wiley Environmental Molecular Sciences Laboratory, Pacific Northwest National Laboratory, Richland, WA 99352.

Fourier transform infrared spectra including resolved spectral doublets for many transitions are reported for gaseous tropolone(OH), tropolone(OD), and (18-O,18-O)tropolone(OH) in the 800 to 150cm^{−1} spectral range. The known zeropoint tunneling splitting values Δ_{0} = 0.974cm^{−1} for tropolone(OH) (Tanaka et al.) and 0.051cm^{−1} for tropolone(OD) (Keske et al.) allow vibrational state-specific tunneling splittings Δ_{v} to be estimated for fundamentals including the three with the strongest OO stretching displacements [cf. for tropolone(OH) ν_{13} = 435.22cm^{−1} with Δ_{13} = 1.71cm^{−1} = 1.76 Δ_{0}, and for tropolone(OD) ν_{13} = 429.631cm^{−1} with Δ_{13} = 0.311cm^{−1} = 6.10 Δ_{0}]. In addition to revealing isotope-dependent Δ_{v} splittings the FTIR study demonstrates effects due to dynamic couplings and potential function anharmonicity. Vibrational reassignments are fostered and the pattern of Δ_{v} tunneling splitting values is interpreted to underscore the intrinsic multidimensionality of the tautomerization process in this nonrigid 15 atom molecule.